inverse of the function f(x) = 3x + 2. State the answer in slope-intercept form.
Answer by jim_thompson5910(35256)

f(x) = 3x + 2

y = 3x + 2

x = 3y + 2

x - 2 = 3y

x/3 - 2/3 = y

y = x/3 - 2/3

y = (1/3)x - 2/3

So the inverse function is